OpenStudy (anonymous):

solve for y' : xy' + y = 1 + y'

OpenStudy (anonymous):

xy' +y = 1 + y' xy' - y' = 1 - y y'(x-1) = 1 - y y' = (1-y)/(x-1)

OpenStudy (anonymous):

thank you so very much!:) i really appreciate it:)

OpenStudy (anonymous):

you're welcome